OpenStudy (anonymous):

PLEASES HELP NOW!! Subtract 5 1/4 - 3 5/8 2 1/8 2 5/8 1 1/8 1 5/8

OpenStudy (anonymous):

plz help now

OpenStudy (anonymous):

\[\left(5+\frac{1}{4}\right)-\left(3+\frac{5}{8}\right)=\left(5+\frac{2}{8}\right)-\left(3+\frac{5}{8}\right)=\frac{42}{8}-\frac{29}{8}=\frac{13}{8}=1+\frac{5}{8} \]

OpenStudy (anonymous):

i have to submit this plz like now

OpenStudy (anonymous):

so wait what is the answer? @robtobey

OpenStudy (anonymous):

1 5/8

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

i have another question can u plz help me

OpenStudy (anonymous):

Xavier's dining room table was unstable because one leg was shorter than the others. He placed a 3/8 inch thick piece of wood under the short leg, and it became even with the other legs, which are 19 1/4 inches long. How many inches long was the short leg before Xavier added the wood? 19 5/8 inches 18 3/4 inches 19/ 7/8 inches 18 7/8 inches